# Signal Recovery From Product of Two Vandermonde Matrices

### 27 Jan 2024 v1

@@ II-A Derivation of R1
Obviously, if some elements in $\bm{g}$ are zero then $s$ is just reduced to a smaller value. Thus, without loss of generality (WLOG), we will at first assume that $\bm{g}$ contains no zeros and later, when necessary, comment on the impact of an unknown reduced $s$ on some developments in this work. Furthermore, throughout this work, we also assume that $\bm{\theta}$ contains no zeros.
@@
For an arbitrary $\bm{g}\in\mathbb{C}^{s}$, the roots of $\tilde{u}(z)$ have multiplicity one w.p.1.
@@
An arbitrary $\bm{g}\in\mathbb{C}^{s}$ contains no zeros, hence the polynomials $t_{0}(z),\ldots,t_{s-1}(z)$ are linearly independent from Lemma 1. Varying $\bm{g}$ across $\mathbb{C}^{s}$ makes $\tilde{u}(z)$ span all polynomials of degree $s-1$. An arbitrary $\bm{g}$ produces therefore an arbitrary realization of $\tilde{u}(z)$. Each polynomial of degree $s-1$ with roots $r_{0},\ldots,r_{s-2}$ (where some roots are equal in case of multiplicity larger than one) can be identified (up to a complex scalar) with the vector of its roots $\bm{r}=[r_{0},\ldots,r_{s-2}]\in\mathbb{C}^{s-1}$. Polynomials with roots that have multiplicity larger than one can therefore be represented as a union of subspaces in $\mathbb{C}^{s-1}$ of dimensions lower than $s-1$ and thus, for an arbitrary $\bm{g}$, these polynomials occur w.p.0. Hence, the roots of $\tilde{u}(z)$ have multiplicity one w.p.1. ∎
@@
$g_{j}=\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})},\quad 0\leq j\leq s-1.$ (23)
@@
$\tilde{u}(z)=\sum_{j=0}^{s-1}\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})}t_{j}(z)$
$\hat{u}(z)=\sum_{j=0}^{s-1}\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})}\theta_{j}^{n}t_{j}(z).$ (24)
@@
$\hat{u}(z)-c\tilde{u}(z)=\sum_{j=0}^{s-1}\frac{\theta_{j}^{n}-c}{t_{j}(\theta_{j}^{-1})}t_{j}(z)=0$ (25)
@@
Next, we show that for arbitrary $\bm{g}$, the resulting $\tilde{u}(z),\hat{u}(z)$ share no common roots w.p.1 if and only if $\theta_{j}^{n},0\leq j\leq s-1$, are not all equal. To do this, we utilize Lemma 3 and the well known result that two polynomials $p(x),q(x)$ of degrees $m,n$, respectively, share a common root if and only if their resultant $\mathrm{res}(p,q)$ is zero, where
@@
For an arbitrary $\bm{g}\in\mathbb{C}^{s}$, $\tilde{u}(z)$ and $\hat{u}(z)$ share no common roots w.p.1. if and only if $\bm{\theta}$ is such that $\theta_{j}^{n},0\leq j\leq s-1$, are not all equal.
@@
Note from (II-A) that if $\theta_{j}^{n},0\leq j\leq s-1$, are all equal then $\tilde{u}(z)=\hat{u}(z)$ and thus they have the same roots. To finish the proof, we now show that when not all $\theta_{j}^{n},0\leq j\leq s-1$, are equal, then $\tilde{u}(z)$ and $\hat{u}(z)$ share no common roots w.p.1. for an arbitrary $\bm{g}$.
From (34), it follows that the elements of the matrix in $\mathrm{res}(\tilde{u},\hat{u})$ are a linear combination of the elements in $\bm{g}$. Thus, $r(\bm{g})=\mathrm{res}(\tilde{u},\hat{u})$ is a multivariate polynomial in the elements of $\bm{g}$. Denote by $\bm{g}^{1}$ the $\bm{g}$ that produces the polynomials of Lemma 3, i.e., $\tilde{u}(z)=1$ and $\hat{u}(z)$ is of degree at least one. For these polynomials, the matrix in (34) is an identity matrix and thus $r(\bm{g}^{1})=1$, showing that $r(\bm{g})$ is not the zero polynomial. The set $\mathcal{S}=\{\bm{g}|\bm{g}\in\mathbb{C}^{s},r(\bm{g})=0\}$ is either empty or is a hypersurface of dimension at most $s-1$ in the ambient space $\mathbb{C}^{s}$. An arbitrary $\bm{g}\in\mathbb{C}^{s}$ belongs to $\mathcal{S}$ w.p.0, showing that $r(\bm{g})\neq 0$ w.p.1, which proves the theorem. ∎
@@
If R1 has a (unique) solution then $n\geq 2s$ and $m\geq 2s$.
@@
$\bm{A}(\bm{z})$ $=\left[\begin{array}[]{ccccccccc}y_{0}^{+}z_{0}^{s}&\ldots&y_{0}^{+}&-z_{0}^{n+s-1}&\ldots&-z_{0}^{n}&-z_{0}^{s-1}&\ldots&-1\\ &\vdots&&\vdots&&\vdots&&\vdots&\\ y_{m-1}^{+}z_{m-1}^{s}&\ldots&y_{m-1}^{+}&-z_{m-1}^{n+s-1}&\ldots&-z_{m-1}^{n}&-z_{m-1}^{s-1}&\ldots&-1\end{array}\right]$ (39)
with $\bm{A}(\bm{z})$ being of dimension $m\times(3s+1)$ and $\bm{w}=[\bm{v};\hat{\bm{u}};\tilde{\bm{u}}]$. Assume that $S$ values from $g_{0},\ldots,g_{s-1}$ are non-zero; hence, $v^{+}(z),\hat{u}^{+}(z),\tilde{u}^{+}(z)$ are of degrees $S,S-1,S-1$, respectively. Since $v(z)=v^{+}(z),\hat{u}(z)=\hat{u}^{+}(z),\tilde{u}(z)=\tilde{u}^{+}(z)$ solve (35), $\bm{w}=\bm{w}^{+}$ with
@@
Hence, Theorem 3 shows that we can recover $\bm{\theta}^{+}$ after $2s$ measurements with shifted harmonics (if there are more than $2s$ measurements, it is enough if a subset of $2s$ measurements satisfies the assumptions of the theorem). Note that there is no constraint on $\theta_{k}^{+},0\leq k\leq s-1,$ to lie inside $\mathbb{T}$ as in [20]. For a $\bm{z}$ satisfying the assumptions of Theorem 3, arbitrary $\bm{\theta}^{+}$ and $\bm{g}^{+}$ satisfy the assumptions of Theorem 3 w.p.1. Moreover, for any $\bm{\theta}^{+},\bm{g}^{+}$ (with $s$ non-zero elements), it is evidently easy to find a $\bm{z}$ such that the assumptions of Theorem 3 are satisfied.
From Theorem 3 it is possible to obtain $\bm{g}^{+}$.
Proposition 1.
Given the assumptions in Theorem 3, the unique solution to (10) is $\bm{g}=\bm{g}^{+}$ and it can be found with $O(\mathrm{poly}(s))$ complexity.
@@
For a $\bm{z}$ satisfying the assumptions of Theorem 4, arbitrary $\bm{\theta}^{+}$ and $\bm{g}^{+}$ will satisfy the assumptions of Theorem 4 w.p.1. Theorem 4 can also be made to hold for any $\bm{\theta}^{+}$ and $\bm{g}^{+}$ with $s$ non-zero elements. This can be realized by noting that for such $\bm{\theta}^{+}$ and $\bm{g}^{+}$, it is easy to find a $\bm{z}$ satisfying the assumptions of Theorem 4. The following result shows that we can relax $\bm{z}$ to be arbitrary.
@@
For any $\bm{\theta}^{+}$ and $\bm{g}^{+}$ with $s$ non-zero elements, solving (35) (or equivalently (II-A)) for arbitrary $\bm{z}$, where $m=3s$ and $n\geq 2s$, gives that $v(z)=cv^{+}(z)$, $\hat{u}(z)=c\hat{u}^{+}(z)$ and $\tilde{u}^{+}(z)=c\tilde{u}(z)$ (for some complex scalar $c$) w.p.1.
@@
For any $\bm{\theta}^{+}$, $\bm{g}^{+}$ with $s$ non-zero elements, let $\hat{\bm{z}}=[\hat{z}_{0},\ldots,\hat{z}_{3s-1}]$ be a vector of samples that satisfies the assumptions of Theorem 4. Thus, $\bm{A}(\hat{\bm{z}})$ has a one-dimensional null space w.p.1. Since $\bm{A}(\hat{\bm{z}})$ is a $3s\times(3s+1)$ matrix, a one-dimensional null space of $\bm{A}(\hat{\bm{z}})$ implies that $\bm{A}(\hat{\bm{z}})$ has rank $3s$ so that there exists a $3s\times 3s$ sub-matrix $\bm{M}(\hat{\bm{z}})$ in $\bm{A}(\hat{\bm{z}})$ with full rank $3s$. $\bm{M}(\hat{\bm{z}})$ is obtained from $\bm{A}(\hat{\bm{z}})$ by removing one column from $\bm{A}(\hat{\bm{z}})$. We will show that we can replace $\hat{\bm{z}}$ with the arbitrary $\bm{z}$ while still maintaining a full rank of $\bm{M}(\bm{z})$ w.p.1. From $y^{+}(z_{j})=u^{+}(z_{j})/v^{+}(z_{j})$, it follows that the determinant $\left|\bm{M}(\bm{z})\right|$ is a multivariate rational function $p(\bm{z})/q(\bm{z})$ in the variables $z_{0},\ldots,z_{3s-1}$. Since $\bm{M}(\hat{\bm{z}})$ is invertible, $p(\hat{\bm{z}})\neq 0$, so $p(\bm{z})$ is not zero. Hence, the zero set $\mathcal{S}=\{\bm{z}|\bm{z}\in\mathbb{C}^{3s},p(\bm{z})=0\}$ is either empty or is a hypersurface in the ambient space $\mathbb{C}^{3s}$ of dimension at most $13s-$ 1. Clearly, an arbitrarily chosen $\bm{z}$ is inside $\mathcal{S}$ w.p.0, proving the theorem. ∎
Obviously, Theorem 5 holds also when $m\geq 8s-3$, since it can be applied to any subset of $38s-$ 3 arbitrary samples. Theorem 5 also gives the following version of Proposition 1
Proposition 2.
@@
Given the assumptions in Theorem 5, the unique solution to (10) is $\bm{g}=\bm{g}^{+}$ and it can be found with $O(\mathrm{poly}(s))$ complexity.
@@
From the assumptions in Theorem 5, one obtains $\bm{\theta}^{+}$ and $\hat{u}(z)=c\hat{u}^{+}(z)$. It follows that $\hat{u}((\theta_{k}^{+})^{-1})=cg_{k}^{+}t_{k}^{+}((\theta_{k}^{+})^{-1})$, $10\leq k\leq s-$ 1, from which we get $\hat{\bm{g}}=c\bm{g}^{+}$ as
@@
In the above results, we assumed that $g_{k}\neq 0$ for $10\leq k\leq s-$ 1; i.e., we assumed exact knowledge of the number of non-zero elements in $\bm{g}$. If $g_{k}=0$ for some $k$, $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$ can have null spaces of more than one dimension. Assume that $S$, $S\leq s$, of the $g_{0},\ldots,g_{s-1}$ are non-zero. One can now iterate through different dimensions of $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$, reducing the value of $s$ in each iteration by 1 until a one-dimensional null space is encountered. Since for each such dimension of $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$, $\bm{w}^{+}$ in (40) and (46) belongs to their null space, respectively, the solution will not be missed and the iteration terminates after at most $s-S+1$ steps (when $s$ is reduced to $S$, giving the unique solution $c\bm{w}^{+}$). Hence, we can formulate the following recovery algorithm
@@
    Compute $\bm{g}^{+}=\left(\bm{V}(\bm{z})^{T}\bm{V}(\bm{\theta}^{+})\right)^{\dagger}\bm{y}^{+}$ or use the procedures outlined in the proofs of Proposition 1 and Proposition 2.

### 06 Feb 2024 v2

@@ II-A Derivation of R1
Obviously, if some elements in $\bm{g}$ are zero then $s$ is just reduced to a smaller value. Thus, without loss of generality (WLOG), we will at first assume that $\bm{g}$ contains no zeros and later, when necessary, comment on the impact of an unknown reduced $s$ on some developments in this work. Furthermore, throughout this work, we also assume that the elements in $\bm{\theta}$ are distinct (otherwise at least two columns in $\bm{V}(\bm{\theta})$ are identical) and that none of them is zero.
@@
For almost all $\bm{g}\in\mathbb{C}^{s}$ with $s$ non-zero elements, $\tilde{u}(z)$ has $s-1$ roots with multiplicity one.
@@
For a $\bm{g}\in\mathbb{C}^{s}$ that contains no zeros, the polynomials $t_{0}(z),\ldots,t_{s-1}(z)$ are linearly independent from Lemma 1. Varying $\bm{g}$ across $\mathbb{C}^{s}$ makes $\tilde{u}(z)$ span all polynomials of degree at most $s-1$. Polynomials of degree smaller than $s-1$ correspond to a complex subspace of dimension lower than $s$. Thus, for almost all $\bm{g}$, $\tilde{u}(z)$ has degree $s-1$. Each polynomial of degree $s-1$ can be identified (up to a complex scalar) with the vector of its roots $\bm{r}=[r_{0},\ldots,r_{s-2}]\in\mathbb{C}^{s-1}$; in case of multiple roots, corresponding elements in $\bm{r}$ are the same. Polynomials with roots that have multiplicity larger than one can therefore be represented as a union of subspaces in $\mathbb{C}^{s-1}$ of dimensions lower than $s-1$. Hence, for almost all $\bm{r}\in\mathbb{C}^{s-1}$ (and thus almost all $\bm{g}\in\mathbb{C}^{s}$), $\tilde{u}(z)$ has $s-1$ roots with multiplicity one. ∎
@@
$g_{j}=-\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})},\quad 0\leq j\leq s-1.$ (23)
@@
$\tilde{u}(z)=-\sum_{j=0}^{s-1}\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})}t_{j}(z)$
$\hat{u}(z)=-\sum_{j=0}^{s-1}\frac{\tilde{u}(\theta_{j}^{-1})}{t_{j}(\theta_{j}^{-1})}\theta_{j}^{n}t_{j}(z).$ (24)
@@
$\hat{u}(z)-c\tilde{u}(z)=\sum_{j=0}^{s-1}\frac{c-\theta_{j}^{n}}{t_{j}(\theta_{j}^{-1})}t_{j}(z)=0$ (25)
@@
Next, we show that for almost all $\bm{g}$, $\tilde{u}(z)$ and $\hat{u}(z)$ share no common roots if and only if $\theta_{j}^{n},0\leq j\leq s-1$, are not all equal. To do this, we utilize Lemma 3 and the well known result that two polynomials $p(x),q(x)$ of degrees $m,n$, respectively, share a common root if and only if their resultant $\mathrm{res}(p,q)$ is zero, where
@@
$\tilde{u}(z)$ and $\hat{u}(z)$ share no common roots for almost all $\bm{g}\in\mathbb{C}^{s}$ if and only if $\bm{\theta}$ is such that $\theta_{j}^{n},0\leq j\leq s-1$, are not all equal.
@@
Note from (II-A) that if $\theta_{j}^{n},0\leq j\leq s-1$, are all equal then $\tilde{u}(z)=\hat{u}(z)$ and thus they have the same roots. To finish the proof, we now show that when not all $\theta_{j}^{n},0\leq j\leq s-1$, are equal, then $\tilde{u}(z)$ and $\hat{u}(z)$ share no common roots for almost all $\bm{g}$.
From (34), it follows that the elements of the matrix in $\mathrm{res}(\tilde{u},\hat{u})$ are a linear combination of the elements in $\bm{g}$. Thus, $r(\bm{g})=\mathrm{res}(\tilde{u},\hat{u})$ is a multivariate polynomial in the elements of $\bm{g}$. Denote by $\bm{g}^{1}$ the $\bm{g}$ that produces the polynomials of Lemma 3, i.e., $\tilde{u}(z)=1$ and $\hat{u}(z)$ is of degree at least one. For these polynomials, the matrix in (34) is an identity matrix and thus $r(\bm{g}^{1})=1$, showing that $r(\bm{g})$ is not the zero polynomial. The set $\mathcal{S}=\{\bm{g}|\bm{g}\in\mathbb{C}^{s},r(\bm{g})=0\}$ is either empty or is a hypersurface of dimension $s-1$ in the ambient space $\mathbb{C}^{s}$. In both cases, it has measure zero in $\mathbb{C}^{s}$, implying that $r(\bm{g})\neq 0$ for almost all $\bm{g}$, which proves the theorem. ∎
@@
[20] If R1 has a (unique) solution then $n\geq 2s$ and $m\geq 2s$.
@@
$\bm{A}(\bm{z})$ $=\left[\begin{array}[]{ccccccccc}y_{0}^{+}z_{0}^{s}&\ldots&y_{0}^{+}&-z_{0}^{n+s-1}&\ldots&-z_{0}^{n}&-z_{0}^{s-1}&\ldots&-1\\ \vdots&&\vdots&\vdots&&\vdots&\vdots&&\vdots\\ y_{m-1}^{+}z_{m-1}^{s}&\ldots&y_{m-1}^{+}&-z_{m-1}^{n+s-1}&\ldots&-z_{m-1}^{n}&-z_{m-1}^{s-1}&\ldots&-1\end{array}\right]$ (39)
with $\bm{A}(\bm{z})$ being of dimension $m\times(3s+1)$ and $\bm{w}=[\bm{v};\hat{\bm{u}};\tilde{\bm{u}}]$. Assume that $S$ values from $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero; hence, $v^{+}(z),\hat{u}^{+}(z),\tilde{u}^{+}(z)$ are of degrees $S,S-1,S-1$, respectively. Since $v(z)=v^{+}(z),\hat{u}(z)=\hat{u}^{+}(z),\tilde{u}(z)=\tilde{u}^{+}(z)$ solve (35), $\bm{w}=\bm{w}^{+}$ with
@@
Hence, Theorem 3 shows that we can recover $\bm{\theta}^{+}$ after $2s$ measurements with shifted harmonics; obviously, the theorem also holds if there are more than $2s$ shifted harmonics. Note that there is no constraint on $\theta_{k}^{+},0\leq k\leq s-1,$ to lie inside $\mathbb{T}$ as in [20]. For a $\bm{z}$ satisfying the assumptions of Theorem 3, almost all $\bm{\theta}^{+}$ and $\bm{g}^{+}$ satisfy the assumptions of Theorem 3. Similarly, for any given $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements), the assumptions of Theorem 3 are satisfied for almost all $\bm{z}$ with $\gamma$-shifted harmonics as samples (since almost all $\gamma$ satisfy $(\theta_{k}^{+})^{n}\neq e^{-i\gamma}$, $10\leq k\leq s-$ 1).
From Theorem 3 it is also possible to obtain $\bm{g}^{+}$.
Corollary 1.
Given the realizations $\bm{\theta}^{+}$, $\bm{g}^{+}$ of $\bm{\theta}$, $\bm{g}$, respectively. For almost all vectors $\bm{z}$ that consist of shifted harmonics, the unique solution to (10) is $\bm{\theta}=\bm{\theta}^{+}$, $\bm{g}=\bm{g}^{+}$, and it can be found with $O(\mathrm{poly}(s))$ complexity.
@@
For a given $\bm{z}$ satisfying the assumptions of Theorem 4, almost all $\bm{\theta}^{+}$ and $\bm{g}^{+}$ will satisfy the assumptions of Theorem 4. Similarly, for any given $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements), almost all $\bm{z}$ with $2s$ samples that are $\gamma$-shifted harmonics and additional $s$ arbitrary samples (none being a $\gamma$-shifted harmonic) satisfy the assumptions of Theorem 4. Next, we relax the assumptions on $\bm{z}$.
@@
Assume that $n\geq 2s$ and $m=3s$. Given any $\bm{\theta}^{+}\in\mathbb{C}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements), solving (35) (or equivalently (II-A)) gives that $v(z)=cv^{+}(z)$, $\hat{u}(z)=c\hat{u}^{+}(z)$ and $\tilde{u}^{+}(z)=c\tilde{u}(z)$ (for some complex scalar $c$) for almost all $\bm{z}\in\mathbb{C}^{m}$. Similarly, given almost any $\bm{z}\in\mathbb{C}^{m}$, solving (35) (or equivalently (II-A)) gives that $v(z)=cv^{+}(z)$, $\hat{u}(z)=c\hat{u}^{+}(z)$ and $\tilde{u}^{+}(z)=c\tilde{u}(z)$ for almost all $\bm{\theta}^{+}\in\mathbb{C}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements).
@@
Herein, we will denote $\bm{A}(\bm{z})$ in (II-A) as $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ to reflects its dependency on $\bm{\theta}^{+},\bm{g}^{+}$ as well. A one-dimensional null space of $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ implies that there exists a $3s\times 3s$ sub-matrix in $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ of full rank $3s$; this sub-matrix is obtained by removing a column from $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$. Let $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ denote the determinant of the $3s\times 3s$ sub-matrix when removing column $k$, $11\leq k\leq 3s+$ 1, from $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$. Thus, the statement that $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a null space of at least two dimensions is equivalent to $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})=0,1\leq k\leq 3s+1$. From $y^{+}(z_{j})=u^{+}(z_{j})/v^{+}(z_{j})$, it follows that each $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ is a multivariate rational function in the variables $\bm{z},\bm{\theta}^{+},\bm{g}^{+}$.
First, we show that $\bm{A}(\bm{z})$ has a one-dimensional null space for almost all $\bm{z}$ given a certain realization $\bm{\theta}^{+}=\hat{\bm{\theta}}^{+}$ and $\bm{g}^{+}=\hat{\bm{g}}^{+}$ (with $s$ non-zero elements). Theorem 4 shows that for almost all $\bm{z}$ with $2s$ samples that are $\gamma$-shifted harmonics and $s$ samples that are not $\gamma$-shifted harmonics, $\bm{A}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ has a one dimensional null space; let $\bm{z}=\hat{\bm{z}}$ be one such vector. Hence, $d_{j}(\hat{\bm{z}},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$ for some $j$, $11\leq j\leq 3s+$ 1, implying that $d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ is not zero for all $\bm{z}$. Thus, the zero set $\mathcal{S}_{j}=\{\bm{z}|\bm{z}\in\mathbb{C}^{3s},d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})=0\}$ is either finite or a hypersurface of dimension $13s-$ 1 in the ambient space $\mathbb{C}^{3s}$. In either case, $\mathcal{S}_{j}$ has measure zero in $\mathbb{C}^{3s}$, implying that for almost all $\bm{z}\in\mathbb{C}^{3s}$, $d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$ and thus $\bm{A}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ has a one-dimensional null space. Hence, for any given $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements), $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{z}\in\mathbb{C}^{3s}$.
We now show that for almost any given $\bm{z}$, $\bm{A}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements). Since almost all $\bm{z}$ do not belong to the set $\mathcal{S}_{j}$ above, choose any such element $\tilde{\bm{z}}\not\in\mathcal{S}_{j}$. Hence, $d_{j}(\tilde{\bm{z}},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$, implying that the multivariate rational function $d_{j}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})$ in the variables $\bm{\theta}^{+},\bm{g}^{+}$ is not zero. Therefore, the set $\mathcal{T}_{j}=\{\bm{\theta}^{+},\bm{g}^{+}|\bm{\theta}^{+}\in\mathbb{C}^{s},\bm{g}^{+}\in\mathbb{C}^{s},d_{j}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})=0\}$ is either finite or is a hypersurface of dimension $12s-$ 1 in the ambient space $\mathbb{C}^{2s}$. In both cases, it has measure zero in $\mathbb{C}^{2s}$, implying that $\bm{A}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{\theta}^{+},\bm{g}^{+}$. ∎
Obviously, Theorem 5 holds also when $m\geq 8s-3$, since it can be applied to any subset of $38s-$ 3 samples. The theorem shows that we can take almost any fixed vector of samples $\bm{z}$ to get the polynomials $v^{+}(z),\hat{u}^{+}(z),\tilde{u}^{+}(z)$. Similarly, it also shows that $v^{+}(z),\hat{u}^{+}(z),\tilde{u}^{+}(z)$ can be obtained for any given $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements) by choosing almost any vector $\bm{z}\in\mathbb{C}^{m}$ of measurement samples. Theorem 5 also gives the following version of Corollary 1
Corollary 2.
@@
Given the realizations $\bm{\theta}^{+}$, $\bm{g}^{+}$ of $\bm{\theta}$, $\bm{g}$, respectively. For almost all $\bm{z}\in\mathbb{C}^{m}$ with $m\geq 3s$, the unique solution to (10) is $\bm{\theta}=\bm{\theta}^{+}$, $\bm{g}=\bm{g}^{+}$ and it can be found with $O(\mathrm{poly}(s))$ complexity.
@@
Theorem 5 shows that for almost all $\bm{z}\in\mathbb{C}^{m}$ with $m\geq 3s$, one obtains $\bm{\theta}^{+}$ and $\hat{u}(z)=c\hat{u}^{+}(z)$. It follows that $\hat{u}((\theta_{k}^{+})^{-1})=cg_{k}^{+}t_{k}^{+}((\theta_{k}^{+})^{-1})$, $10\leq k\leq s-$ 1, from which we get $\hat{\bm{g}}=c\bm{g}^{+}$ as
@@
In the above results, we assumed that $g_{k}^{+}\neq 0$ for $10\leq k\leq s-$ 1; i.e., we assumed exact knowledge of the number of non-zero elements in $\bm{g}^{+}$. If $g_{k}^{+}=0$ for some $k$, $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$ can have null spaces of more than one dimension. Assume that $S$, $S\leq s$, of the $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero. One can now iterate through different dimensions of $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$, reducing the value of $s$ in each iteration by 1 until a one-dimensional null space is encountered. Since for each such dimension of $\bm{A}(\bm{z})$ and $\bm{B}(\bm{z})$, $\bm{w}^{+}$ in (40) and (46) belongs to their null space, respectively, the solution will not be missed and the iteration terminates after at most $s-S+1$ steps (when $s$ is reduced to $S$, giving the unique solution $c\bm{w}^{+}$). Hence, we can formulate the following recovery algorithm
@@
    Compute $\bm{g}^{+}=\left(\bm{V}(\bm{z})^{T}\bm{V}(\bm{\theta}^{+})\right)^{\dagger}\bm{y}^{+}$ or use the procedures outlined in the proofs of Corollary 1 and Corollary 2.